# Gauging tensor networks with belief propagation

### v4

@@ 2.1 Tensor network states and the Vidal gauge
where $*{}^{*}$ denotes complex conjugation and the right-hand side of Eq. (2) represents the identity matrix (provided there is an appropriate normalization on $\Gamma_{v}$ and $\Lambda_{e}$). The open indices in Eq. (2) are directed along the edge between $v$ and one of its $d$ neighbors $v_{i}$. The Vidal gauge can be defined for an arbitrary (finite or infinite) TNS and the bond tensors can be used to provide a rank-one approximation [^1] Methods like boundary MPS [117, 120, 155] or corner transfer matrix renormalization group (CTMRG) [156, 121, 157, 158, 159] exist for obtaining higher-rank approximations of environments in loopy tensor networks, though they have primarily been applied to tensor networks on grids or other regular lattices. Other approximate tensor network contraction methods include a variety of generalizations of BP [154, 143, 150, 147, 152, 160, 161, 162], tensor renormalization methods [163, 164, 165], and a growing list of other general approximate contraction algorithms [166, 167, 29]. of the environment in a general loopy network — allowing rapid, approximate local updates to be performed on the TNS. When the TNS has a tree structure, the Vidal gauge allows one to define an exact reduced, orthogonal basis at any site in the network. We will introduce a new method for transforming a TNS into the Vidal gauge using belief propagation, independent of the network structure. The essential steps are outlined in Fig. 3 and in the following we will provide the explicit details.
@@ 2.4 Using belief propagation to gauge a tensor network state
The belief propagation (BP) gauging routine Start with a TNS consisting of site tensors $\{T_{v}\}$. Form the closed network corresponding to the square of the norm of the TNS. Initialize the belief propagation message tensors $\{M_{e},M_{\bar{e}}\}$ of the normed network to arbitrary, positive definite matrices [^4] While initializing to positive definite matrices may not be strictly necessary it guarantees the message tensors are positive semidefinite at each step and therefore their square root can be taken.. Perform iterations of belief propagation, where each iteration corresponds to an update of each message tensor via Eq. (5), up to some stopping criteria. Gauge the TNS with the resulting message tensors via Eqs. (21), (22), and (23). This yields the TNS in the Vidal gauge with site tensors $\{\Gamma_{v}\}$ and bond tensors $\{\Lambda_{e}\}$.
@@ 3.1 Quantifying the distance to the Vidal gauge of a tensor network state
Equation (29) essentially measures the average degree to which Eqs. (1) and (2) are satisfied by the tensors in the TNS. Here, $||A-B||_{1^{\prime}}=\left|\left|\frac{A}{{\rm Tr}(A)}-\frac{B}{{\rm Tr}(B)}\right|\right|_{1}$ and $||M||_{1}$ represents the trace norm of a matrix $M$. Both matrices in Eq. (29) are scaled to trace unity in order to remove any dependence of $\mathcal{C}$ on the normalization of the network.
@@
$O\left(\mathcal{C}^{(n)}\right)\approx\frac{1}{2|E|}\sum_{v}\sum_{i=1}^{|N(v)|}\left|\left|M^{(n-1)}_{v,v_{i}}-M^{(n)}_{v,v_{i}}\right|\right|_{1^{\prime}},$ (31)
where $n$ is the current iteration of BP. This can be derived using Eqs. (5), (21), (22), and (23) and comparing to Eq. (29) [^5] The following reasonable assumptions are necessary in this derivation: $O(\mathcal{C}^{(n)})\approx O(\mathcal{C}^{(n+1)})$ and $O\left(\left|\left|\frac{A}{{\rm Tr}(A)}-\frac{B}{{\rm Tr}(B)}\right|\right|_{1}\right)\approx O\left(\left|\left|\frac{B^{-1/2}AB^{-1/2}}{{\rm Tr}(B^{-1/2}AB^{-1/2})}-\mathbb{I}\right|\right|_{1}\right)$. This allow us to use BP to gauge a TNS to a prescribed order of magnitude of $\mathcal{C}$ without any significant computational overhead from checking the value of $\mathcal{C}$.
@@ 3.2 Accelerating tensor network gauging with belief propagation
Figure 4: a-b) Time to reach $\mathcal{C}=10^{-10}$ for a random TNS — with tensor elements drawn from the normal distribution of mean 00 and standard deviation $1$ — of bond dimension $\chi$ on an $L\times L$ square lattice. a) Scaling with $\chi^{5}$ for $L=12$. The plot ranges to $\chi=32$ where we see that eager gauging and simple update gauging take $1.82$ and $2.81$ times longer than belief propagation gauging respectively. b) Scaling with $L^{2}$ for $\chi=15$. The plot ranges to $L=35$ where we see that eager gauging and simple update gauging take $1.75$ and $2.63$ times longer than belief propagation gauging respectively. c) Distance to the Vidal gauge $\mathcal{C}$ versus walltime for gauging a random tensor network state with elements drawn from the normal distribution of mean 00 and standard deviation $1$. The bond dimension of the TNS is $\chi=20$ and two lattices are shown: the hexagonal TNS for $10$ rows and $10$ columns of hexagons, and a $200$ site random regular graph of fixed degree $3$. In plots a-c) different line styles correspond to different gauging methods (with the legend at the top of the figure) and all walltimes are in seconds. Iterations of each algorithm are performed using a fixed Suzuki-Trotter decomposition-inspired schedule of updates over the edges of the TNS (see main text for more details). d) Number of iterations required to reach $\mathcal{C}=10^{-8}$ for the TNS corresponding to the square root of the classical Ising partition function (which has bond dimension $2$) versus inverse temperature $\beta$. A longitudinal field of $h=0.5$ is included. Results are for a $20\times 20\times 20$ cubic lattice. Different colors correspond to different update schedules for the message tensors in a given iteration - with further details in the main text - and the total walltime is directly proportional to the number of iterations.
@@ 3.3 Approximate gate application with belief propagation
Figure 5: a) Simple update gate evolution in the Vidal gauge. i) - ii) The bond tensors and gate $O_{v,w}$ (marked by the red dotted line) are combined into a single composite tensor $\Theta_{v,w}$. ii) - iii) $\Theta_{v,w}$ is SVDd and singular values of the resulting bond tensor $\tilde{\Lambda}_{v,w}$ can be discarded in order to truncate the bond down to a desired dimension $\chi^{\prime}$. iii) - iv) Resolutions of identity, using the original bond tensors, are inserted on the exposed edges. iv) - v) The inverse bond tensors are absorbed, resulting in the updated site tensors $\tilde{\Gamma}_{v}$ and $\tilde{\Gamma}_{w}$ and the bond tensor $\tilde{\Lambda}_{v,w}$. b) The analog of simple update gate evolution using BP message tensors, equivalent to ‘full update’ gate evolution where the environment is products of message tensors. The steps are analogous to the simple update procedure for the Vidal gauge. The new message tensors on the bond $v,w$ is taken to be $\tilde{M}_{v,w}=\tilde{M}_{w,v}=\Lambda$, the diagonal tensor from the SVD whose square root is also absorbed by $U_{v}$ and $V_{w}$. a) and b) are equivalent updates of the state and represent ideal rank-one gate applications if, before applying the gates, either the state in a) is regauged using one of the Vidal gauging methods or BP is run on the state in b) to get updated fixed point message tensors. More efficient versions of these procedures (with and without the gate) are pictured in Figs. 10 and 11.
@@ 3.6 Gauging infinite tensor networks with belief propagation
Figure 8: Gauging infinite tensor networks with BP gauging. i) Infinite 2D TNS (PEPS) with a $2\times 2$ unit cell. Unit cells are separated by red dotted lines. ii) Unit cell of the infinite 2D TNS but with additional periodic edges added (dashed lines). iii) Upon gauging the periodic unit cell new site tensors and bond tensors are determined which put it into the Vidal gauge. The bond tensors are not necessarily identical, and can be unique for each edge (in this example there are $8$). iv) The resulting site tensors and bond tensors from the periodic TNS determine the Vidal gauge for the infinite lattice. These tensors can be used to directly extract observables from the infinite TNS, for example via Eq. (36).
@@ Appendix A Simple update
Figure 10: ‘Reduced tensor’ version of simple update described in Fig. 5. The example pictured shows a two-site gate $O_{v,w}$ being applied to neighbouring sites $v$ and $w$ with coordination numbers of $3$ and $4$ respectively. The dimension of the bond along the edge $e=(v,w)$ is $\chi$. i) - ii) The bond tensors are absorbed onto the site tensors. ii) - iii) A QR decomposition is performed on both the left and right site tensors. iii) - iv) The resulting $R_{v}$ and $R_{w}$ tensors, the bond tensor $\Lambda_{v,w}$, and the gate are combined together to form the composite tensor $\Theta_{v,w}$. iv) - v) The composite tensor is SVDd (with the bond being truncated down to dimension $\chi^{\prime}$) and the resulting singular values become the new bond tensor $\tilde{\Lambda}_{v,w}$. v) - vi) The $U_{v}$ and $V_{w}$ matrices from the SVD are multiplied back with $Q_{v}$ and $Q_{w}$. vi) - vii) Resolutions of identity using the previous bond tensors are inserted. vii) - viii) The inverses are absorbed to yield the updated bond tensors and site tensors. Such a procedure could be generalized to long-range gates by matrix product operator decompositions of the gate.
Figure 11: Simple update step when applying an identity matrix. This step is the workhorse of simple update gauging. The example pictured is two neighbouring sites with co-ordination numbers of $3$ and $4$ respectively. The bond dimension along the edge $e=(v,w)$ is $\chi$. i) - ii) The bond tensors are absorbed onto the site tensors. ii) - iii) A QR decomposition is performed on both the left and right site tensors. iii) - iv) The resulting $R$ tensors are combined together to form the composite $\Theta_{v,w}$. iv) - v) The composite tensor is SVDd (with the bond being truncated down to dimension $\chi^{\prime}\leq\chi$) and the resulting singular values become the new bond tensor $\tilde{\Lambda}_{v,w}$. v) - vi) The $U_{v}$ and $V_{w}$ matrices from the SVD are multiplied back with $Q_{v}$ and $Q_{w}$. vi) - vii) Resolutions of identity using the previous bond tensors are inserted. vii) - viii) The inverses are absorbed to yield the updated bond tensors and site tensors.

### v5

@@ 2.1 Tensor network states and the Vidal gauge
where denotes complex conjugation and the right-hand side of Eq. (2) represents the identity matrix (provided there is an appropriate normalization on $\Gamma_{v}$ and $\Lambda_{e}$). The open indices in Eq. (2) are directed along the edge between $v$ and one of its $d$ neighbors $v_{i}$. The Vidal gauge can be defined for an arbitrary (finite or infinite) TNS and the bond tensors can be used to provide a rank-one approximation [^1] Methods like boundary MPS [117, 120, 155] or corner transfer matrix renormalization group (CTMRG) [156, 121, 157, 158, 159] exist for obtaining higher-rank approximations of environments in loopy tensor networks, though they have primarily been applied to tensor networks on grids or other regular lattices. Other approximate tensor network contraction methods include a variety of generalizations of BP [154, 143, 150, 147, 152, 160, 161, 162], tensor renormalization methods [163, 164, 165], and a growing list of other general approximate contraction algorithms [166, 167, 29]. of the environment in a general loopy network — allowing rapid, approximate local updates to be performed on the TNS. When the TNS has a tree structure, the Vidal gauge allows one to define an exact reduced, orthogonal basis at any site in the network. We will introduce a new method for transforming a TNS into the Vidal gauge using belief propagation, independent of the network structure. The essential steps are outlined in Fig. 3 and in the following we will provide the explicit details.
@@ 2.4 Using belief propagation to gauge a tensor network state
The belief propagation (BP) gauging routine
  • Start with a TNS consisting of site tensors $\{T_{v}\}$. Form the closed network corresponding to the square of the norm of the TNS.
  • Initialize the belief propagation message tensors $\{M_{e},M_{\bar{e}}\}$ of the normed network to arbitrary, positive definite matrices [^4] While initializing to positive definite matrices may not be strictly necessary it guarantees the message tensors are positive semidefinite at each step and therefore their square root can be taken..
  • Perform iterations of belief propagation, where each iteration corresponds to an update of each message tensor via Eq. (5), up to some stopping criteria.
  • Gauge the TNS with the resulting message tensors via Eqs. (21), (22), and (23). This yields the TNS in the Vidal gauge with site tensors $\{\Gamma_{v}\}$ and bond tensors $\{\Lambda_{e}\}$.
@@ 3.1 Quantifying the distance to the Vidal gauge of a tensor network state
[…]
@@
[…]
@@ 3.2 Accelerating tensor network gauging with belief propagation
Figure 4: a-b) Time to reach $\mathcal{C}=10^{-10}$ for a random TNS — with tensor elements drawn from the normal distribution of mean 0 and standard deviation 11 — of bond dimension $\chi$ on an $L\times L$ square lattice. a) Scaling with $\chi^{5}$ for $L=12$. The plot ranges to $\chi=32$ where we see that eager gauging and simple update gauging take $1.82$ and $2.81$ times longer than belief propagation gauging respectively. b) Scaling with $L^{2}$ for $\chi=15$. The plot ranges to $L=35$ where we see that eager gauging and simple update gauging take $1.75$ and $2.63$ times longer than belief propagation gauging respectively. c) Distance to the Vidal gauge $\mathcal{C}$ versus walltime for gauging a random tensor network state with elements drawn from the normal distribution of mean 0 and standard deviation 11. The bond dimension of the TNS is $\chi=20$ and two lattices are shown: the hexagonal TNS for $10$ rows and $10$ columns of hexagons, and a $200$ site random regular graph of fixed degree $3$. In plots a-c) different line styles correspond to different gauging methods (with the legend at the top of the figure) and all walltimes are in seconds. Iterations of each algorithm are performed using a fixed Suzuki-Trotter decomposition-inspired schedule of updates over the edges of the TNS (see main text for more details). d) Number of iterations required to reach $\mathcal{C}=10^{-8}$ for the TNS corresponding to the square root of the classical Ising partition function (which has bond dimension 22) versus inverse temperature $\beta$. A longitudinal field of $h=0.5$ is included. Results are for a $20\times 20\times 20$ cubic lattice. Different colors correspond to different update schedules for the message tensors in a given iteration - with further details in the main text - and the total walltime is directly proportional to the number of iterations.
@@ 3.3 Approximate gate application with belief propagation
Figure 5: a) Simple update gate evolution in the Vidal gauge. i) - ii) The bond tensors and gate $O_{v,w}$ (marked by the red dotted line) are combined into a single composite tensor $\Theta_{v,w}$. ii) - iii) $\Theta_{v,w}$ is SVDd and singular values of the resulting bond tensor $\tilde{\Lambda}_{v,w}$ can be discarded in order to truncate the bond down to a desired dimension $\chi^{\prime}$. iii) - iv) Resolutions of identity, using the original bond tensors, are inserted on the exposed edges. iv) - v) The inverse bond tensors are absorbed, resulting in the updated site tensors $\tilde{\Gamma}_{v}$ and $\tilde{\Gamma}_{w}$ and the bond tensor $\tilde{\Lambda}_{v,w}$. b) The analog of simple update gate evolution using BP message tensors, equivalent to ‘full update’ gate evolution where the environment is products of message tensors. The steps are analogous to the simple update procedure for the Vidal gauge. The new message tensors on the bond $v,w$ is taken to be $\tilde{M}_{v,w}=\tilde{M}_{w,v}=\Lambda$, the diagonal tensor from the SVD whose square root is also absorbed by $U_{v}$ and $V_{w}$. a) and b) are equivalent updates of the state and represent ideal rank-one gate applications if, before applying the gates, either the state in a) is regauged using one of the Vidal gauging methods or BP is run on the state in b) to get updated fixed point message tensors. More efficient versions of these procedures (with and without the gate) are pictured in Figs. 10 and $1$.
@@ 3.6 Gauging infinite tensor networks with belief propagation
Figure 8: Gauging infinite tensor networks with BP gauging. i) Infinite 2D TNS (PEPS) with a $2\times 2$ unit cell. Unit cells are separated by red dotted lines. ii) Unit cell of the infinite 2D TNS but with additional periodic edges added (dashed lines). iii) Upon gauging the periodic unit cell new site tensors and bond tensors are determined which put it into the Vidal gauge. The bond tensors are not necessarily identical, and can be unique for each edge (in this example there are 88). iv) The resulting site tensors and bond tensors from the periodic TNS determine the Vidal gauge for the infinite lattice. These tensors can be used to directly extract observables from the infinite TNS, for example via Eq. (36).
@@ Appendix A Simple update
Figure 10: ‘Reduced tensor’ version of simple update described in Fig. 5. The example pictured shows a two-site gate $O_{v,w}$ being applied to neighbouring sites $v$ and $w$ with coordination numbers of $3$ and 44 respectively. The dimension of the bond along the edge $e=(v,w)$ is $\chi$. i) - ii) The bond tensors are absorbed onto the site tensors. ii) - iii) A QR decomposition is performed on both the left and right site tensors. iii) - iv) The resulting $R_{v}$ and $R_{w}$ tensors, the bond tensor $\Lambda_{v,w}$, and the gate are combined together to form the composite tensor $\Theta_{v,w}$. iv) - v) The composite tensor is SVDd (with the bond being truncated down to dimension $\chi^{\prime}$) and the resulting singular values become the new bond tensor $\tilde{\Lambda}_{v,w}$. v) - vi) The $U_{v}$ and $V_{w}$ matrices from the SVD are multiplied back with $Q_{v}$ and $Q_{w}$. vi) - vii) Resolutions of identity using the previous bond tensors are inserted. vii) - viii) The inverses are absorbed to yield the updated bond tensors and site tensors. Such a procedure could be generalized to long-range gates by matrix product operator decompositions of the gate.
Figure 11: Simple update step when applying an identity matrix. This step is the workhorse of simple update gauging. The example pictured is two neighbouring sites with co-ordination numbers of $3$ and 44 respectively. The bond dimension along the edge $e=(v,w)$ is $\chi$. i) - ii) The bond tensors are absorbed onto the site tensors. ii) - iii) A QR decomposition is performed on both the left and right site tensors. iii) - iv) The resulting $R$ tensors are combined together to form the composite $\Theta_{v,w}$. iv) - v) The composite tensor is SVDd (with the bond being truncated down to dimension $\chi^{\prime}\leq\chi$) and the resulting singular values become the new bond tensor $\tilde{\Lambda}_{v,w}$. v) - vi) The $U_{v}$ and $V_{w}$ matrices from the SVD are multiplied back with $Q_{v}$ and $Q_{w}$. vi) - vii) Resolutions of identity using the previous bond tensors are inserted. vii) - viii) The inverses are absorbed to yield the updated bond tensors and site tensors.